### v1

@@ 2 Longest Paths and Cycles on the Real Line
First we prove the sufficiency of (i) by contradiction. Let $C$ be a longest cycle on $P$, and orient its edges to obtain a directed cycle. Suppose, for the sake of contradiction, that the edge $(p_{a},p_{b})$ of $C$ does not intersect the median. We may assume w.l.o.g. that $p_{a},p_{b}<0$. The sum of vertex degrees strictly on the left and right side of the median are the same, and the edges that contain 00 in their interior contribute 1 to both sums. Consequently, $C$ contains an edge $(p_{c},p_{d})$ with $p_{c},p_{d}>0$; or (when $n$ is odd) there are two edges incident to the median, say $(p_{c},0)$ and $(0,p_{d})$ with $p_{c},p_{d}>0$. In the first case, we can replace edges $(p_{a},p_{b})$ and $(p_{c},p_{d})$ with $(p_{a},p_{c})$ and $(p_{b},p_{d})$. In the second case, replace $(p_{a},p_{b})$ and $(p_{c},0)$ with $(p_{a},p_{c})$ and $(p_{b},0)$. In both cases, we obtain a longer (undirected) spanning cycle, contradicting the maximality of $C$.
The necessity of (i) can be proved by a counting argument similar to that of Lemma 2-(i).
Now, we prove (ii) by contradiction. Without loss of generality, let $0\in P$ be the median of $P$. Suppose that the median is incident to two edges $(p_{c},0)$ and $(0,p_{d})$ with $p_{c},p_{d}>0$. Then, there is a point in $P$ to the right of 0 incident to an edge of $C$ that does not contain 0 in its interior. Denote this edge by $(p_{a},p_{b})$, where $p_{a},p_{b}<0$. We can replace edges $(p_{a},p_{b})$ and $(p_{c},0)$ with $(p_{a},p_{c})$ and $(p_{b},0)$ to obtain a longer spanning cycle, contradicting the maximality of $C$.
@@
To prove the first part of (iii), note that if $n=2k{+}1$, then the median is the $(k{+}1)$-st point of $P$, that we denote by $p_{0}$. Let $C$ be a longest cycle on $P$. It is implied from (i) and (ii) that exactly one edge of $C$ (which is incident to $p_{0}$) does not contain $I$. The remaining $n{-}1=2k$ edges contain $I$.
@@
For the second claim in (iii), let $C$ be a spanning cycle on $P$ in which fewer than $2k$ edges contain $I$. Orient the edges of $C$ to obtain a directed cycle. The sum of degrees of the leftmost $k{+}1$ (resp., rightmost $k$) vertices is $22k{+}$ 2 (resp., $2k$), and the edges containing $I$ have fewer than $2k$ left (resp., right) endpoints. Consequently, the leftmost $k{+}1$ (resp., rightmost $k$) points in $P$ induce at least two edges (resp., one edge) of $C$. Therefore, $C$ contains two edges, $(p_{a},p_{b})$ and $(p_{c},p_{d})$, such that $p_{a},p_{b}$ are to the left of $I$ and $p_{b},p_{d}$ are to the right of $I$. We can replace these two edges with $(p_{a},p_{c})$ and $(p_{b},p_{d})$, to obtain a spanning cycle $C^{\prime}$ that traverses $I$ two more times than $C$. In particular, we have $|C^{\prime}|\geq|C|+2\,|I|=|C|+2h$, hence $|C|\leq|C^{\prime}|-2h\leq|C_{\max}|-2h$, where $C_{\max}$ is a longest cycle on $P$. ∎
@@ 3 Noncrossing Longest Paths
Let $n\geq 1$ be an integer. In this section, we construct $n$ points for which the longest spanning path is unique and noncrossing. This can be easily observed for $n<5$: For example, for $n=4$, any spanning path of the vertices of a triangle and a point in the interior is noncrossing. Thus, we will now assume that $n\geq 5$. In Section 2, we uncovered some structural properties of longest paths for $n$ points on a line. Here we show how to construct a 2-dimensional point set starting with $n$ points on the $x$-axis and then assigning $y$-coordinates to the points. We show that the longest path is unique and noncrossing. We describe our construction for the case where $n$ is even; the construction for the case where $n$ is odd follows with some minor changes. The following theorem summarizes our result in this section.
@@
In Section 3.1 we give an overview of our construction for an even number of points. The details and proofs are given in Section 3.2. The case of odd paths is considered in Section 3.3.
3.1 A path with an even number of points: An overview
For $k\geq 3$, consider a set $P$ of $n=2k$ points $p_{i}$ on the $x$-axis such that $p_{1}=(0,0)$ and $p_{i}=(i,0)$ for $i=-1,\pm 2,\dots,\pm k$, as illustrated in Figure 2(a). Our construction would work even if we set $p_{1}=(1,0)$; however, for a reason that will become clear in Section 3.3, we set $p_{1}$ differently. The longest spanning path for this point set is not unique. In fact, Lemma 2 implies that any spanning path with endpoints $p_{1}$ and $p_{-1}$ and with all edges crossing the $y$-axis is a longest path. Conversely, any longest path must have endpoints $p_{1}$ and $p_{-1}$, and its edges must cross the $y$-axis. Let $\cal H$ be the set of these paths. Let $P^{\prime}$ be the point set obtained by assigning to each point $p_{i}$ a $y$-coordinate $y_{i}$ such that, as illustrated in Figure 2(b), the following holds:
$\frac{1}{8k}=y_{1}\gg y_{-2}\gg y_{2}\gg y_{-3}\gg y_{3}\gg\cdots\gg y_{-k}\gg y_{k}\gg y_{-1}=0.$
The value $y_{1}$ is much larger than $y_{-2}$, which is in turn much larger than $y_{2}$ and so on. Notice that the largest $y$-coordinate $y_{1}$ is $1/8k$ which is much smaller than $1$. Due to the small $y$-coordinates, a longest path $H^{\prime}$ on $P^{\prime}$ corresponds to a path $H\in\cal H$. The length of $H^{\prime}$ is roughly the length of $H$ plus a very small value $\Delta(H^{\prime})$, which depends on the new $y$-coordinates. Let $e_{1}$ be the only edge of $H^{\prime}$ incident to $p_{1}$. Since $p_{1}$ has a very large $y$-coordinate compared to other points, the contribution of $e_{1}$ to $\Delta(H^{\prime})$ is larger than the contribution of other edges. The contribution of $e_{1}$ is maximized if it connects $p_{1}$ to the nearest plausible neighbor, which is $p_{-2}$; this can be observed from Figure 2(b). Therefore $e_{1}=p_{1}p_{-2}$. By a similar argument, $p_{-2}$ gets connected to $p_{2}$, and so on. It follows that the path $H^{\prime}$ is unique and it is $p_{1},p_{-2},p_{2},p_{-3},p_{3},\dots,p_{-k},p_{k},p_{-1}.$ This path is $y$-monotone, and hence noncrossing; see Figure 2(b).
Note.  Figures 2(a) and 2(b) are not to scale. The $y$-coordinates should be small enough so that all points lie almost on the $x$-axis (We exaggerated the $y$-coordinates to facilitate readability). Moreover, if we orient the path from $p_{1}$ towards $p_{-1}$, then the extension of every directed edge intersects all edges that follow.
Figure 2: Illustration of the construction of a longest path for $2k$ points. The figure is not to scale as the real $y$-coordinates are very small so that the points lie almost on the $x$-axis. (a) Lifting $p_{1}$ to the $y$-coordinate $1/8k$. (b) The final longest path.
3.2 A path with an even number of points: Details
Recall the set $P$ of $2k$ points, $k\geq 3$, on the $x$-axis, described in the previous section and illustrated in Figure 2(a). We say that an edge $e$ intersects the $y$-axis if the intersection of $e$ and the $y$-axis is not empty (the intersection could be an endpoint of $e$). The longest paths for points on a line were characterized in Lemma 2. Denote by $\cal H$ the set of all longest spanning paths on $P$.
@@ 3.2 A path with an even number of points: Details
Let $0\leq\varepsilon\leq\frac{1}{8k}$ be a real number. Suppose that every point in $P$ is perturbed by a distance of at most $\varepsilon$. Let $P^{\prime}$ be the new point set after perturbation. Then, the order of the points along any longest path for $P^{\prime}$ is the same as the order of the points along some path in $\cal H$.
@@
The length of any path on $P$ is an integer. Therefore, any path in $\cal H$ is at least 1 unit longer than any path not in $\cal H$.
Let $H^{\prime}$ be any longest path on $P^{\prime}$. The difference between its length and the length of any path in $\cal H$ is at most $(2k{-}1)\cdot 2\varepsilon$ because $H^{\prime}$ has $12k{-}$ 1 edges, each edge has $2$ endpoints, and each endpoint is at distance at most $\varepsilon$ from its corresponding point in $P$. Since $\varepsilon\leq\frac{1}{8k}$ the difference is less than $1/2$. Therefore, $H^{\prime}$ cannot correspond to a path that is not in $\cal H$, so $H^{\prime}$ corresponds to a path in $\cal H$ with the same order of points. ∎
Our plan is to assign new $y$-coordinates to the points of $P$ to obtain a point set $P^{\prime}$ for which the longest path is $y$-monotone and unique. The new $y$-coordinates will be at most $\frac{1}{8k}$, and thus, by Lemma 4, the longest path $H^{\prime}$ of $P^{\prime}$ will correspond to a path in $\cal H$. We will make $H^{\prime}$ correspond to the path $p_{1},p_{-2},p_{2},p_{-3},p_{3},\allowbreak\dots,p_{k},p_{-1}$, which is in $\cal{H}$ (by Lemma 2) and depicted in Figure 2(b). We assign to each point $p_{i}$ the $y$-coordinate $y_{i}$ such that the following holds:
$y_{1}\gg y_{-2}\gg y_{2}\gg y_{-3}\gg y_{3}\gg\cdots\gg y_{-k}\gg y_{k}\gg y_{-1}.$
We set $y_{1}=\frac{1}{8k}$, $y_{-1}=0$, and use the following lemma to identify the remaining $y$-coordinates.
@@
There exists a real number $\delta$, with $0<\delta<y_{1}$, such that if $0\leq y_{i}\leq\delta$ for each $i\neq 1$ then the longest path on $P^{\prime}$ connects $p_{1}$ to $p_{-2}$.
@@
Since each $y_{i}$ is at most $1/8k$, Lemma 4 implies that any longest path $H^{\prime}$ on $P^{\prime}$ corresponds to a path $H$ in $\cal H$. Due to small $y$-coordinates, we have $|H^{\prime}|=|H|+\Delta(H^{\prime})$ for some small value $\Delta(H^{\prime})\geq 0$ which depends on the new $y$-coordinates. Specifically, we have
$|H^{\prime}|$ $=\sum_{(p_{i},p_{j})\in E(H^{\prime})}|p_{i}p_{j}|=\sum_{(p_{i},p_{j})\in E(H^{\prime})}\sqrt{|i-j|^{2}+|y_{i}-y_{j}|^{2}}$
$=|H|+\sum_{(p_{i},p_{j})\in E(H^{\prime})}\left(\sqrt{|i-j|^{2}+|y_{i}-y_{j}|^{2}}-|i-j|\right)=|H|+\Delta(H^{\prime}),$
where $0\leq\sqrt{|i-j|^{2}+|y_{i}-y_{j}|^{2}}-|i-j|\leq|y_{i}-y_{j}|\leq\max\{y_{i},y_{j}\}$.
Recall from Lemma 2 that $p_{1}$ is an endpoint of any longest path in $\cal H$. Moreover, $p_{1}$ is connected to a point (different from $p_{-1}$) to the left of the $y$-axis. For $j\in\{2,\dots,k\}$ let $\ell_{j}$ be the Euclidean distance between $p_{1}$ and the point $(-j,0)$, and let $\Delta_{j}$ be the difference of their $x$-coordinates as in Figure 2(a). The contribution of $p_{1}p_{-j}$ to $|H^{\prime}|$ would be at least $\ell_{j}-\delta$ (when $p_{-j}$ has $y$-coordinate $\delta$) and at most $\ell_{j}$ (when $p_{-j}$ has $y$-coordinate 00). The contribution of the corresponding edge to $|H|$ would be $\Delta_{j}$. Hence the contribution of $p_{1}p_{-j}$ to $\Delta(H^{\prime})$ would be at least $\ell_{j}-\delta-\Delta_{j}$ and at most $\ell_{j}-\Delta_{j}$. An easy calculation shows that
$\ell_{2}-\Delta_{2}>\ell_{3}-\Delta_{3}>\dots>\ell_{k}-\Delta_{k};$
this is also implied by the fact that $\Delta_{i+1}{-}\Delta_{i}=1$ while $\ell_{i+1}{-}\ell_{i}<1$. If we set $\delta<(\ell_{2}-\Delta_{2})-(\ell_{3}-\Delta_{3})$, then the contribution of $p_{1}p_{-2}$ to $\Delta(H^{\prime})$ is at least
$\ell_{2}-\delta-\Delta_{2}>\ell_{2}-\Delta_{2}-((\ell_{2}-\Delta_{2})-(\ell_{3}-\Delta_{3}))=\ell_{3}-\Delta_{3},$
which is larger than the contribution of any other plausible edge $p_{1}p_{-j}$. Since the $y$-coordinates of all other points are less than $\delta$, any other edge of $H^{\prime}$ contributes less than $\delta$ to $\Delta(H^{\prime})$. By setting
$\delta=\frac{(\ell_{2}-\Delta_{2})-(\ell_{3}-\Delta_{3})}{2k-1},$
the contribution of $p_{1}p_{-2}$ exceeds the sum of the contributions of the remaining $22k-$ 2 edges of $H^{\prime}$. Thus, for this choice of $\delta$ the longest path $H^{\prime}$ connects $p_{1}$ to $p_{-2}$. ∎
By Lemma 5, we have a specific value $\delta$ such that the longest path includes edge $(p_{1},p_{-2})$. Now we set $y_{-2}=\delta$ and repeat the arguments of Lemma 5, with $y_{-2}$ and $p_{-2}$ (instead of $y_{1}$ and $p_{1}$). This implies that the next edge of the longest path will connect $p_{-2}$ to $p_{2}$. Repeating this $52k{-}$ 5 more times, we obtain the unique longest path $p_{1},p_{-2},p_{2},p_{-3},p_{3},\dots,p_{k},p_{-1}$, as in Figure 2(b); in each of the last two steps, there is only one remaining plausible edge (namely, $p_{-k}p_{k}$ from $p_{-k}$, and $p_{k}p_{-1}$ from $p_{k}$). This path is $y$-monotone and hence is noncrossing.

### v2

@@ 2 Longest Paths and Cycles on the Real Line
First we prove the necessity of (i) by contradiction. Let $C$ be a longest cycle on $P$, and orient its edges to obtain a directed cycle. Suppose, for the sake of contradiction, that the edge $(p_{a},p_{b})$ of $C$ does not intersect the median. We may assume without loss of generality that $p_{a},p_{b}<0$. The sum of vertex degrees strictly on the left and right side of the median are the same, and the edges that contain 0 in their interior contribute 1 to both sums. Consequently, either $C$ contains an edge $(p_{c},p_{d})$ with $p_{c},p_{d}>0$; or (when $n$ is odd) the two edges incident to the median, say $(p_{c},0)$ and $(0,p_{d})$ satisfy $p_{c},p_{d}>0$. In the first case, we can replace the undirected edges $p_{a}p_{b}$ and $p_{c}p_{d}$ with $p_{a}p_{c}$ and $p_{b}p_{d}$. In the second case, replace $p_{a}p_{b}$ and $p_{c}0$ with $p_{a}p_{c}$ and $p_{b}0$. In both cases, we obtain a longer (undirected) spanning cycle, contradicting the maximality of $C$.
The sufficiency of (i) can be proved by a counting argument similar to that of Lemma 3.
Next, we prove (ii) by contradiction. Suppose that the median, $0\in P$, is incident to the edges $(p_{c},0)$ and $(0,p_{d})$ with $p_{c},p_{d}>0$. As the sum of vertex degrees strictly on the left and right side of the median are the same, there is an edge of $C$ strictly to the left of 0; this contradicts part (i).
@@
To prove the first claim in (iii), note that if n=2k+1n=2k{+}1, then the median is the (k+1)(k{+}1)-st point in $P$, that we denote by $p_{0}$. Let $C$ be a longest cycle on $P$. Now (i) and (ii) imply that exactly one edge of $C$ (which is incident to $p_{0}$) does not contain $I$. The remaining n1=2kn{-}1=2k edges contain $I$.
@@
For the second claim in (iii), let $C$ be a spanning cycle on $P$ in which fewer than $2k$ edges contain $I$. Orient the edges of $C$ to obtain a directed cycle. The sum of degrees of the leftmost k+1k{+}1 ( respectively, rightmost $k$) vertices is 2k+22k{+}2 ( respectively, $2k$), and the edges containing $I$ have fewer than $2k$ left ( respectively, right) endpoints. Consequently, the leftmost k+1k{+}1 ( respectively, rightmost $k$) points in $P$ induce at least two edges ( respectively, one edge) of $C$. Therefore, $C$ contains two edges, $(p_{a},p_{b})$ and $(p_{c},p_{d})$, such that $p_{a},p_{b}$ are to the left of $I$ and $p_{c},p_{d}$ are to the right of $I$. We can replace these two edges with $(p_{a},p_{c})$ and $(p_{b},p_{d})$, to obtain a spanning cycle $C^{\prime}$ that traverses $I$ two more times than $C$. In particular, we have $|C^{\prime}|\geq|C|+2\,|I|=|C|+2h$, hence $|C|\leq|C^{\prime}|-2h\leq|C_{\max}|-2h$, where $C_{\max}$ is a longest cycle on $P$. ∎
@@ 3 Noncrossing Longest Paths
Let $n\geq 1$ be an integer. In this section, we construct $n$ points for which the longest spanning path is unique and noncrossing. This can be easily observed for $n<5$: For example, for $n=4$, every spanning path of the vertices of a triangle and a point in the interior is noncrossing. Thus, we will now assume that $n\geq 5$. In Section 2, we uncovered some structural properties of longest paths for $n$ points on a line. Here we show how to construct a 2-dimensional point set starting with $n$ points on the $x$-axis and then assigning $y$-coordinates to the points. We show that the longest path is unique and noncrossing. We describe our construction for the case where $n$ is even; the construction for the case where $n$ is odd follows with some minor changes. The following theorem summarizes the main result in this section.
@@
In Section 3.1 we prove a generalization of Theorem 1. The case of odd paths is handled in Section 3.2.
3.1 A Path with an Even Number of Points
In this section, we prove a stronger statement than Theorem 1 for even $n$, and show that a set $P$ of an even number of points on the $x$-axis can be perturbed, by slightly increasing their $y$-coordinates, to a point set $P^{\prime}$ for which the longest spanning path is unique and $y$-monotone (hence noncrossing).
@@ 3.1 A Path with an Even Number of Points
For every even integer $n\geq 2$ and every set $P$ of $n$ real numbers, there exists a set $P^{\prime}$ of $n$ points in the plane with the following properties:
  1. 1.
    the $x$-projection of $P^{\prime}$ is $P$;
  2. 2.
    the $x$-projection of every longest path on $P^{\prime}$ is a longest path on $P$;
  3. 3.
    the longest spanning path on $P^{\prime}$ is unique and noncrossing.
The rest of this subsection is dedicated to the proof of Lemma 6. The statement is trivial for $n=2$, so we may assume $n\geq 4$. Let $P=\{x_{-1},x_{1},x_{-2},x_{2},\ldots,x_{-k},x_{k}\}$ for some integer $k=n/2$ such that $x_{i}<x_{j}$ if and only if $i<j$. Furthermore, we may assume without loss of generality that the median of $P$ is 0. Since $n$ is even, then $0\notin P$ and we have $x_{i}<0$ if and only of $i<0$.
We construct the point set $P^{\prime}=\{p_{-1},p_{1},p_{-2},p_{2},\ldots,p_{-k},p_{k}\}$ where $p_{i}=(x_{i},y_{i})$ for some $y_{i}\geq 0$ to be specified later. In particular, the $y$-coordinates will satisfy
$y_{1}>y_{-2}>y_{2}>y_{-3}>y_{3}>\cdots>y_{-k}>y_{k}>y_{-1}=0.$
First we choose $y_{1}$. For every $Q\subseteq P$ with $|Q|\geq 3$, let $\beta(Q)$ be the difference between the length of the longest and the second longest spanning paths on $Q$; and let $\beta={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min}_{Q\subseteq P}\beta(Q)$. We set $y_{1}=\frac{\beta}{8k}$. Since $P$ is a set of integers, $\beta\geq 1$. Then Corollary 1 implies that the $x$-projection of every longest path on $P^{\prime}$ is a longest path on $P$. We choose the remaining $y$-coordinates in k1k-1 iterations and maintain the following invariant for $i\in\{1,\ldots,k\}$.
Invariant ($M_{i}$).
There exist real numbers $\frac{\beta}{8k}=y_{1}>y_{-2}>y_{2}>\cdots>y_{-i}>y_{i}>y_{-1}=0$ such that for each $i\in\{1,\ldots,k\}$, every longest spanning path on $P^{\prime}$ starts with $p_{1},p_{-2},p_{2},\dots,p_{-i},p_{i}$ as a subpath.
The Invariant $(M_{1})$ vacuously holds as the 1-vertex path $p_{1}$ is a subpath of every spanning path on $P^{\prime}$.
Figure 2: Illustration of the construction of a longest path for $2k$ points. The figure is not true to scale as the real $y$-coordinates are very small so that the points lie almost on the $x$-axis. (a) Lifting $(x_{1},0)$ to the $y$-coordinate $\frac{\beta}{8k}$. (b) The final longest path.
Note.  Figures 2(a) and 2(b) are not true to scale. The $y$-coordinates should be small enough so that all points lie almost on the $x$-axis (we exaggerated the $y$-coordinates to facilitate readability).
Overview.
The rough idea of the remaining proof is as follows. Due to the small $y$-coordinates, a longest path $H^{\prime}$ on $P^{\prime}$ corresponds to a longest path $H$ on $P$. The length of $H^{\prime}$ is roughly the length of $H$ plus a very small value $Y(H^{\prime})$, which depends on the new $y$-coordinates. Let $e_{1}$ be the only edge of $H^{\prime}$ incident to $p_{1}$. Since $p_{1}$ has a very large $y$-coordinate compared to other points, the contribution of $e_{1}$ to $Y(H^{\prime})$ is larger than the contribution of other edges. The contribution of $e_{1}$ is maximized if it connects $p_{1}$ to the nearest plausible neighbor, which is $p_{-2}$; this can be observed from Figure 2(b). Therefore, $e_{1}=p_{1}p_{-2}$. By a similar argument, $p_{-2}$ gets connected to $p_{2}$, and so on. It follows that the path $H^{\prime}$ is unique and it is $p_{1},p_{-2},p_{2},p_{-3},p_{3},\dots,p_{-k},p_{k},p_{-1}.$ This path is $y$-monotone, and hence noncrossing; see Figure 2(b). We continue with details on how to choose the remaining $y$-coordinates.
Lemma 7.
If Invariant $(M_{1})$ holds, then we can choose $y_{-2}$ and $y_{2}$ such that $y_{1}>y_{-2}>y_{2}>0$ and Invariant $(M_{2})$ holds.
@@
Since Invariant $(M_{1})$ holds, every longest spanning path of $P^{\prime}$ starts with $p_{1}$. We choose $y_{-2}$ and $y_{2}$ such that $y_{1}>y_{-2}>y_{2}>0$, and that every longest spanning path of $P^{\prime}$ starts with the subpath $p_{1},p_{-2},p_{2}$.
Let $H^{\prime}$ be a longest spanning path on $P^{\prime}$. Notice that $H^{\prime}$ corresponds to a path $H$ on $P$; denote by $\mathcal{H}$ the set of all longest spanning paths on $P$. By the triangle inequality and our constraints on the $y$-coordinates, we have $|H^{\prime}|=|H|+Y(H^{\prime})$ for some $Y(H^{\prime})\geq 0$ that depends on the new $y$-coordinates. In particular, $Y(H^{\prime})\leq|P|\cdot\frac{\beta}{8k}$.
By Lemma 3, $x_{1}$ is an endpoint of every path in $\mathcal{H}$. By Lemma 2(ii), the edge of $H$ incident to $x_{1}$ intersects the median of $P$, and so this edge is $x_{1}x_{-j}$ for some $j\in\{2,\ldots,k\}$. For $j\in\{2,\dots,k\}$ let $\Delta_{j}=|x_{1}-x_{-j}|$ and let $\ell_{j}$ be the Euclidean distance between points $p_{1}$ and $(x_{-j},0)$ in $\mathbb{R}^{2}$; as in Figure 2(a). The contribution of $p_{1}p_{-j}$ to $|H^{\prime}|$ is $|p_{1}p_{-j}|$ . On the one hand, we have $|p_{1}p_{-j}|\geq\ell_{j}-y_{-j}\geq\ell_{j}-y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}$ by the triangle inequality and the assumption $y_{-2}\geq y_{-j}$. On the other hand, we have $|p_{1}p_{-j}|\leq\ell_{j}$. The contribution of the corresponding edge $x_{1}x_{-j}$ to $|H|$ is exactly $\Delta_{j}$. Hence the contribution of $p_{1}p_{-j}$ to $Y(H^{\prime})$ is at least $\ell_{j}-y_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}}-\Delta_{j}$ and at most $\ell_{j}-\Delta_{j}$. The triangle inequality (for the triangle $p_{1}p_{-j}p_{-(j+1)}$) yields $\ell_{j+1}<\ell_{j}+(\Delta_{j+1}-\Delta_{j})$. Consequently, we have
$\ell_{2}-\Delta_{2}>\ell_{3}-\Delta_{3}>\dots>\ell_{k}-\Delta_{k}.$
If we set $y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}<(\ell_{2}-\Delta_{2})-(\ell_{3}-\Delta_{3})$, then the contribution of $p_{1}p_{-2}$ to $Y(H^{\prime})$ is at least
$\ell_{2}-y_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}}-\Delta_{2}>\ell_{2}-\Delta_{2}-((\ell_{2}-\Delta_{2})-(\ell_{3}-\Delta_{3}))=\ell_{3}-\Delta_{3},$
which is larger than the contribution of every other plausible edge $p_{1}p_{-j}$. Since the $y$-coordinates of all other points are less than $y_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}}$, every other edge of $H^{\prime}$ contributes less than $y_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}}$ to $Y(H^{\prime})$. By setting
$y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}=\frac{(\ell_{2}-y_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}}-\Delta_{2})-(\ell_{3}-\Delta_{3})}{2k-1},$
the contribution of $p_{1}p_{-2}$ exceeds the sum of the contributions of the remaining 2k22k-2 edges of $H^{\prime}$. After rearrangement, we get
$y_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}}=\frac{(\ell_{2}-\Delta_{2})-(\ell_{3}-\Delta_{3})}{2k}.$
Thus, for this choice of $y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}$ the longest path $H^{\prime}$ contains the edge $p_{1}p_{-2}$. Note also that $\ell_{2}<\Delta_{2}+y_{1}$ by the triangle inequality, which gives $y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}<\ell_{2}-\Delta_{2}<y_{1}$, as required.
By Lemma 3, $x_{-2}$ is not an endpoint of HH. [^2] Note that $P\setminus\{x_{1}\}$ contains an odd number of points, but we do not use the characterization of longest paths on odd sets (Lemma 4). In fact, if we remove $x_{1}$ from $H$, we do not necessarily obtain a longest path on $P\setminus\{x_{1}\}$. By Lemma 2(ii), every edge of $H$ intersects the median of $P$. Consequently, $x_{-2}$ is adjacent to $x_{1}$ and some point $x_{j}$, $j\in\{2,\ldots,k\}$. Similarly to the choice of $y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}$, we can choose $y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}$ so that the longest path $H^{\prime}$ contains the edge $p_{-2}p_{2}$ and $0<y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}<y_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}$. This establishes Invariant $(M_{2})$. ∎
@@
For every $i\in\{1,\ldots,k-1\}$, if Invariant $(M_{i})$ holds, then we can choose $y_{-(i+1)}$ and $y_{i+1}$ such that $y_{i}>y_{-(i+1)}>y_{i+1}>0$ and Invariant $(M_{i+1})$ holds.
@@
The case $i=1$ is handled in Lemma 7. In general, assume that Invariant $(M_{i})$ holds for some $i\in\{2,\ldots,k-2\}$. Then the longest spanning path $H^{\prime}$ on $P^{\prime}$ starts with the subpath $p_{1},p_{-2},p_{2},\ldots,p_{-i},p_{i}$. Let $\widehat{P}$ be the set of real numbers $x_{j}\in P$ with $|j|\geq i$; and let $\widehat{P}^{\prime}$ be the corresponding subset of $P^{\prime}$. Note that $\widehat{P}$ contains an even number of points, and the two closest points to its median are $p_{i}$ and $p_{-1}$. Recall that $y_{i}<y_{1}=\frac{\beta}{8k}$. By the choice of $\beta$, Invariant $(M_{1})$ holds for $\widehat{P}$ with the current value of $y_{i}$. In particular, $p_{i}$ and $p_{-1}$ are the endpoints of every longest spanning path on $\widehat{P}^{\prime}$. Lemma 7 yields $y_{-(i+1)}$ and $y_{i+1}$ such that $y_{i}>y_{-(i+1)}>y_{i+1}>0$ and every longest spanning path of $\widehat{P}^{\prime}$ contains the subpath $p_{i},p_{-(i+1)},p_{i+1}$. Consequently, every longest spanning path of $P^{\prime}$ contains the subpath $p_{1},p_{-2},p_{2},\dots,p_{i},p_{-(i+1)},p_{i+1}$.
Finally, for i=k1i=k-1, we can choose $y_{-k}$ and $y_{k}$ arbitrarily so that $y_{k-1}>y_{-k}>y_{k}>y_{-1}=0$. Invariant $(M_{k-1})$ implies that $p_{1}$ and $p_{-1}$ are the endpoints of every longest spanning path $H^{\prime}$ on $P^{\prime}$. Therefore, $p_{k-1}$ is not an endpoint of $H^{\prime}$, but its adjacent edges intersect the $y$-axis by Lemma 2(ii). There is only one remaining plausible edge from $p_{k-1}$ and $p_{-k}$: $p_{k-1}p_{-k}$ and $p_{-k}p_{k}$, respectively. ∎
Proof of Lemma 6.
By Lemma 8, Invariant $(M_{k})$ holds. This implies that there are $y$-coordinates for all points in $P^{\prime}$ such that every longest path $H^{\prime}$ on $P^{\prime}$ corresponds to a longest path on $P$. Furthermore, $H^{\prime}$ contains $p_{1},p_{-2},p_{2},\dots,p_{-k},p_{k}$ as a subpath. By Lemma 2(i), $p_{1}$ and $p_{-1}$ are the endpoints of $H^{\prime}$. It now follows that $H^{\prime}=p_{1},p_{-2},p_{2},\dots,p_{-k},p_{k},p_{-1}$. Since this path is $y$-monotone, it is noncrossing. ∎